OpenStudy (anonymous):

find the length of the curve y=x^2 from x=1 to x=4

OpenStudy (anonymous):

Im really confused on how to even start this problem

OpenStudy (dumbcow):

Arc length of a curve f(x) is given as: \[\int\limits_{}^{}\sqrt{1+f'(x)^{2}}dx\]

myininaya (myininaya):

right!

OpenStudy (dumbcow):

yeah it pretty much the same thing...simplify yours a little more

OpenStudy (anonymous):

agree w/ cow.

myininaya (myininaya):

\[\int\limits_{1}^{4}\sqrt{1+((x^2)')^2} dx=\int\limits_{1}^{4}\sqrt{1+(2x)^2}dx\]

myininaya (myininaya):

\[\int\limits_{1}^{4}\sqrt{1+4x^2} dx\]

OpenStudy (dumbcow):

so f(x) = x^2 limits from 1 to 4 what he did ^^

myininaya (myininaya):

she*

OpenStudy (dumbcow):

haha my bad...my apologies

myininaya (myininaya):

you will need a trig substitution here

myininaya (myininaya):

let tan(theta)=2x

OpenStudy (anonymous):

I'm sorry I am really not understanding

myininaya (myininaya):

all you have to do is use the formula

OpenStudy (anonymous):

really?

OpenStudy (anonymous):

is it just simply as plugging the 1 and 4 in at this point \[\int\limits\limits_{1}^{4}\sqrt{1+4x ^{2}}\]

myininaya (myininaya):

you have to integrate

myininaya (myininaya):

then plug in limits

myininaya (myininaya):

remember the fundamental thm of calculus

myininaya (myininaya):

i thought you are having trouble finding the length so the trouble is you don't know how to integrate?

OpenStudy (anonymous):

ummm ok now i see I will try to find the answer thanks

OpenStudy (anonymous):

integration is easy i just forgot how to handle this problem

myininaya (myininaya):

ok